OpenStudy (anonymous):

If the linear equation 2x = 9 − 3y is written in standard form, what is the value of B? wouldnt it be 9?

OpenStudy (anonymous):

@BTaylor do u kno this?

OpenStudy (btaylor):

when you put it in linear form, you are solving for y. So you'll get: 3y=-2x + 9 y= -2/3 x +3 so b = 3.